### v2

@@ 2.1 ROS 2
Internally, ROS 2 implementation uses a layered architecture depicted in Figure 1. The figure also shows the main differences between C++ and Rust R2R nodes, which will be detailed below.
@@ 2.2.1 Execution model of futures runtime
The futures::join!() macro can be used to group multiple futures and wait for the completion of the group as a whole. The effect of joining the futures is that their code will not be executed in parallel. From the point of view of the executor, the group is treated as one task; if one future gets ready, the entire group becomes ready. When the group starts executing, it will poll all ready tasks in the group. The effect of this grouping in the executor is shown later in an example with ROS in Figure 4.
@@
Tokio is a popular multi-threaded runtime. It seems to be designed to maximize throughput rather than time determinism. Their scheduling policies are complex and difficult to understand due to the use of abstraction layers. The scheduling policies are partially described in the documentation [25] and can be outlined as follows: Each worker thread has a LIFO slot, which is used for dequeuing ready tasks at most three times in a row. We believe, this improves cache locality while also avoiding starvation. Then, tasks are dequeued from the local ready queue, which can hold up to 256 tasks. If the worker cannot dequeue tasks from the local or global (shared) ready queues (they are empty), it steals half of the tasks in the other worker’s local queue. The victim of the theft is chosen as the first worker, with a non-empty queue when iterating workers with a random starting position. Such a behavior is clearly unsuitable for real-time applications. We will evaluate Tokio experimentally in Section 4.
@@ 2.4 Feature comparison of rclcpp, R2R and rclrs
Before looking in detail at the R2R client library and its execution model, we provide a high-level comparison of features implemented in R2R and the other ROS Rust client library rclrs. As both are community-supported, they both lack some features implemented in the C++ client library rclcpp. The comparison of all three libraries is summarized in Table 1 and commented in more detail below.
@@ 2.5 R2R execution model
R2R [5] is one of ROS 2 client libraries for Rust. Its execution model differs from C++ because callback execution is managed by a Rust async runtime rather than by R2R itself as shown in Figure 1. R2R is only responsible for “sampling” the events from ROS entities and pushing them to the async runtime via MPSC channels.
@@
The registration of entity callbacks differs from C++, where methods for creating timers or subscriptions take the callback as a parameter. In R2R, the corresponding methods return the receiving end of the associated MPSC channel, and the callback is bound to it by spawning an async task created from an async block, which periodically awaits the timer expiration or the message from the channel. See Section 2.5 for an example. This way, if a subscription channel contains more than one message, the callback is consecutively executed for all messages available in the channel.
{listing}
Registration of a subscription callback in R2R.{minted}rust let subscription_future = subscription.for_each(|msg| async move // Callback code for msg processing here ); executor.spawn(subscription_future);
{listing}
R2R setup equivalent to one single-threaded executor in C++ {minted}[linenos]rust local_executor.run_until_stalled(); // initialization loop node.spin_once(Duration::seconds(1)); // sampling local_executor.run_until_stalled(); // execution
The execution of the callbacks is carried on by the executors. An example of the main R2R loop with the futures local executor is shown in Section 2.5. The call to run_until_stalled before the loop at line 1 is necessary if we want to avoid memory allocation in the loop. The call initializes the executor by moving the tasks from the incoming vector to the active list, which involves memory allocation. Since the spin_once function was not called before, no tasks were sampled, and no callback code is going to be executed.
@@
Figure 2 shows the above described sequence of operations in time. Besides the R2R thread running the loop from Section 2.5 (denoted as Node S), it shows publisher and DDS threads involved in the process.
@@
Note that the callbacks may not always be executed in the FIFO order as described above. One such case can happen if spin_once is called multiple times without calling run_until_stalled in between. This is demonstrated in the example in Figure 3. The rightmost column shows in which order would be the callbacks called by a call to run_until_stalled following the spin_once calls.
@@ 2 A B C B D A C B1, B2, D, A, C
A similar effect can happen if multiple futures are joined as described in Section 2.2.1. An example of how the callbacks would be executed when grouping them via the join macro is shown in Figure 4.
@@ 2.6 Comparison of C++ and R2R execution models
The main difference between C++ and R2R execution models is summarized in Figure 5. R2R does not sample events in executors because waiting on a wait set is a synchronous blocking operation, which should not be executed from an async task.
However, if the R2R application is organized as in Section 2.5, the execution is equivalent to using a ROS C++ single-threaded executor with a minor difference that the priorities of subscription and timer callbacks are swapped. This means that the response-time analysis proposed in [2] can be applied to R2R as is while taking into account the difference in priorities.
{listing}
R2R sampling loop to mimic ROS C++ multi-threaded executor. {minted}rust loop node.spin_once(Duration::seconds(1));
To achieve the behavior of a C++ multi-threaded executor with R2R, sampling must be performed in a dedicated thread with a loop shown in Section 2.6. The effect of different C++ callback group configurations can be accomplished as follows:
@@ 3 Achieving deterministic real-time operation with R2R
An example of the proposed structure is given in Section 3. The priority of the main thread is set on lines 14–18, before creating ROS context and node at lines 20 and 21. A single subscriber is then created on line 23, followed by associating it with a callback implemented as an async move block (lines 24–26). Line 27 spawns a new OS thread with the given priority. The code executed by the thread is a closure defined at lines 5–10. It creates a new local executor (line 6), spawns to it a new async task from the passed future (line 8) and runs the executor loop (line 9).
@@
Execution of callbacks in lower-priority threads ensures that one can use schedulability analysis, which does not need to take into account ROS specifics and depends only on the scheduling policy of the OS scheduler. In Section 4, we demonstrate this by running each callback in a thread pinned to a single CPU and with priority set according to the rate-monotonic ordering and use uni-processor response-time analysis [1] [^6] Application of more complex analysis techniques should be possible and is left for future work. to predict the response times.
@@
{listing}
Proposed structure of a real-time R2R application {minted}[linenos]rust fn spawn_in_thread(future: impl Future, priority: ThreadPriority) let thread = ThreadBuilder::default() .policy(RealTime(Fifo)) .priority(priority) .spawn(move |_| let mut local_executor = executor::LocalPool::new(); let spawner = local_executor.spawner(); spawner.spawn_local(future).unwrap(); local_executor.run(); );
fn main() -> Result<(), Box<dyn Error>> thread_priority::unix::set_thread_priority_and_policy( thread_priority::thread_native_id(), ThreadPriority::try_from(MAIN_PRIORITY)?, RealTime(Fifo), )?;
let ctx = r2r::Context::create()?; let mut node = r2r::Node::create(ctx, "example", "")?;
let subs = node.subscribe("/topic", QosProfile::default())?; let future = subs.for_each(move |msg: Msg| async move // process msg ); spawn_in_thread(future, ThreadPriority::try_from(CALLBACK_PRIORITY)?);
loop node.spin_once(SPIN_TIMEOUT);
@@ 4 Experimental evaluation
To evaluate the proposed application structure and compare it with other alternatives, we designed a set synthetic benchmarks. These are described in Section 4.1. Evaluation on a more complex autonomous driving application follows in Section 4.2.
@@ 4.1 Synthetic benchmarks
To evaluate the proposed application structure and compare response times obtained with different setups, we develop a benchmark application composed of five topics (see Figure 6). Messages are published periodically to the topics by a publisher node running in a dedicated process. The second node subscribes to the topics and executes callbacks with a specific fixed execution time. The publication periods and callback execution times are given in Table 2 and lead to CPU utilization of 90%.
@@ 4.1.3 Implementation variants of the subscriber node
R2R-based subscriber node with futures local executor running interleaved with the spin_once function in the same thread, as in Section 2.5.
@@
The variant proposed in Section 3, that is R2R-subscriber node with each callback running in a separate thread with real-time priority set according to the rate-monotonic (RM) priority assignment (the smaller period, the higher priority) in a futures local executor. The spin_once function is called in a loop in the main thread with the highest priority.
@@ 4.1.4 Synthetic benchmark results
To investigate the effect of RMW/DDS thread priority, we compare the end-to-end latencies of our proposed futures-rt with and without giving the RMW/DDS threads the highest priority. The results in Figure 8 show that without setting the priority of RMW/DDS threads to be higher than the priorities of the callback threads, the callbacks fail to complete before the deadlines given by the publication periods.
@@ 4.2 Complex autonomous driving application
Our ALKS implementation can drive a real Porsche Cayenne car by connecting to its FlexRay buses and communicating with the onboard control units. However, in this paper, we evaluate the version running against the simulated vehicle in the CARLA simulator [7]. The reason is that this version contains three ROS nodes written using R2R, whereas the real car version has only two R2R-based nodes. The high-level architecture of the application is depicted in Figure 9.
@@
The overview of threads in the nodes and their SCHED_FIFO priorities is given in Table 3. Only the ALKS node runs callbacks in more than one thread.
@@ 4.2.1 Latency evaluation
To show that R2R is suitable for implementing practical real-time applications, we run the above-described ALKS application and trace it with LTTng. As opposed to the synthetic benchmark, here we did not restrict the threads to run on a single CPU core. From the traces, we obtain end-to-end latencies of the odometry chain (marked with red arrows in Figure 9). This chain goes through all three R2R-based nodes in the application. Specifically, it starts at the reception of the IMU message from CARLA in the FlexRay adapter. In the associated callback, we retrieve other information from CARLA using its API and publish the obtained data to the corresponding ROS topic. The ALKS node then receives (among others) the odometry message, converts it from relative to absolute values, and stores it for later use. It also publishes the converted absolute values for use in the FlexRay visualizer. Flexray visualizer just stores the received odometry values in memory for use in other callbacks.
Figure 10 shows histograms obtained from a trace recorded during approximately 3 minutes of execution of the ALKS application. We removed few initial samples from the trace to filter out outliers from the warm-up phase.
The histogram of measured odometry chain end-to-end latencies is shown in Figure 10(a). In about 99% of cases, the latency is below 0.8 ms. The same figure also shows execution time (duration) of the first callback in the chain (IMU). As this callback communicates with the CARLA simulator over the network, its duration and jitter is dominated by this communication latency and this propagates further down the chain. Figure 10(b) shows execution time histograms of other callbacks in the chain. In the ALKS node, the majority of callback executions exhibits execution time jitter of about 0.2 ms, which is caused by waiting for mutexes protecting data shared with the timer callback running in another thread. A few outliers between 0.3 and 2 ms can be caused by the fact that the experiment was executed on a machine connected to the network and running other development tools, which might cause some interference. The jitter of the callback in the FlexRay visualizer node is very low, as the node is single-threaded and does not need to use mutexes in the callbacks.
@@
Figure 10(d) shows the execution time histogram of the timer callback in the ALKS node. The left peak represents the cases where the MPC planner was not invoked, whereas the peak around 8 ms includes the cases with planner invocations. Note that the end-to-end latency of the odometry chain is influenced by the timer callback only thourgh short critical sections for accessing shared data. Long execution time of the MPC planner has no effect on the end-to-end latency of the odometry chain. This demonstrates that R2R with the futures runtime and the structure proposed in this paper is capable of achieving deterministic end-to-end latencies of the chain of callbacks executing in different nodes.

### v3

@@ 2.1 ROS 2
Internally, ROS 2 implementation uses a layered architecture depicted in Figure˜1. The figure also shows the main differences between C++ and Rust R2R nodes, which will be detailed below.
@@ 2.2.1 Execution model of futures runtime
The futures::join!() macro can be used to group multiple futures and wait for the completion of the group as a whole. The effect of joining the futures is that their code will not be executed in parallel. From the point of view of the executor, the group is treated as one task; if one future gets ready, the entire group becomes ready. When the group starts executing, it will poll all ready tasks in the group. The effect of this grouping in the executor is shown later in an example with ROS in Figure˜6.
@@
Tokio is a popular multi-threaded runtime. It seems to be designed to maximize throughput rather than time determinism. Their scheduling policies are complex and difficult to understand due to the use of abstraction layers. The scheduling policies are partially described in the documentation [25] and can be outlined as follows: Each worker thread has a LIFO slot, which is used for dequeuing ready tasks at most three times in a row. We believe, this improves cache locality while also avoiding starvation. Then, tasks are dequeued from the local ready queue, which can hold up to 256 tasks. If the worker cannot dequeue tasks from the local or global (shared) ready queues (they are empty), it steals half of the tasks in the other worker’s local queue. The victim of the theft is chosen as the first worker, with a non-empty queue when iterating workers with a random starting position. Such a behavior is clearly unsuitable for real-time applications. We will evaluate Tokio experimentally in Section˜4.
@@ 2.4 Feature comparison of rclcpp, R2R and rclrs
Before looking in detail at the R2R client library and its execution model, we provide a high-level comparison of features implemented in R2R and the other ROS Rust client library rclrs. As both are community-supported, they both lack some features implemented in the C++ client library rclcpp. The comparison of all three libraries is summarized in Table˜1 and commented in more detail below.
@@ 2.5 R2R execution model
R2R [5] is one of ROS 2 client libraries for Rust. Its execution model differs from C++ because callback execution is managed by a Rust async runtime rather than by R2R itself as shown in Figure˜1. R2R is only responsible for “sampling” the events from ROS entities and pushing them to the async runtime via MPSC channels.
@@
The registration of entity callbacks differs from C++, where methods for creating timers or subscriptions take the callback as a parameter. In R2R, the corresponding methods return the receiving end of the associated MPSC channel, and the callback is bound to it by spawning an async task created from an async block, which periodically awaits the timer expiration or the message from the channel. See Figure˜2 for an example. This way, if a subscription channel contains more than one message, the callback is consecutively executed for all messages available in the channel.
Figure 2: Registration of a subscription callback in R2R.
let subscription_future = subscription.for_each(|msg| async move {
// Callback code for msg processing here
executor.spawn(subscription_future);
Figure 3: R2R setup equivalent to one single-threaded executor in C++
local_executor.run_until_stalled(); // initialization
loop {
node.spin_once(Duration::seconds(1)); // sampling
local_executor.run_until_stalled(); // execution
The execution of the callbacks is carried on by the executors. An example of the main R2R loop with the futures local executor is shown in Figure˜3. The call to run_until_stalled before the loop at line 1 is necessary if we want to avoid memory allocation in the loop. The call initializes the executor by moving the tasks from the incoming vector to the active list, which involves memory allocation. Since the spin_once function was not called before, no tasks were sampled, and no callback code is going to be executed.
@@
Figure˜4 shows the above described sequence of operations in time. Besides the R2R thread running the loop from Figure˜3 (denoted as Node S), it shows publisher and DDS threads involved in the process.
@@
Note that the callbacks may not always be executed in the FIFO order as described above. One such case can happen if spin_once is called multiple times without calling run_until_stalled in between. This is demonstrated in the example in Figure˜5. The rightmost column shows in which order would be the callbacks called by a call to run_until_stalled following the spin_once calls.
@@ 2 A B C B D A C B1, B2, D, A, C
A similar effect can happen if multiple futures are joined as described in Section 2.2.1. An example of how the callbacks would be executed when grouping them via the join macro is shown in Figure˜6.
@@ 2.6 Comparison of C++ and R2R execution models
The main difference between C++ and R2R execution models is summarized in Figure˜7. R2R does not sample events in executors because waiting on a wait set is a synchronous blocking operation, which should not be executed from an async task.
However, if the R2R application is organized as in Figure˜3, the execution is equivalent to using a ROS C++ single-threaded executor with a minor difference that the priorities of subscription and timer callbacks are swapped. This means that the response-time analysis proposed in [2] can be applied to R2R as is while taking into account the difference in priorities.
Figure 8: R2R sampling loop to mimic ROS C++ multi-threaded executor.
loop {
node.spin_once(Duration::seconds(1));
To achieve the behavior of a C++ multi-threaded executor with R2R, sampling must be performed in a dedicated thread with a loop shown in Figure˜8. The effect of different C++ callback group configurations can be accomplished as follows:
@@ 3 Achieving deterministic real-time operation with R2R
An example of the proposed structure is given in Figure˜9. The priority of the main thread is set on lines 14–18, before creating ROS context and node at lines 20 and 21. A single subscriber is then created on line 23, followed by associating it with a callback implemented as an async move block (lines 24–26). Line 27 spawns a new OS thread with the given priority. The code executed by the thread is a closure defined at lines 5–10. It creates a new local executor (line 6), spawns to it a new async task from the passed future (line 8) and runs the executor loop (line 9).
@@
Execution of callbacks in lower-priority threads ensures that one can use schedulability analysis, which does not need to take into account ROS specifics and depends only on the scheduling policy of the OS scheduler. In Section˜4, we demonstrate this by running each callback in a thread pinned to a single CPU and with priority set according to the rate-monotonic ordering and use uni-processor response-time analysis [1] [^6] Application of more complex analysis techniques should be possible and is left for future work. to predict the response times.
@@
Figure 9: Proposed structure of a real-time R2R application
fn spawn_in_thread(future: impl Future, priority: ThreadPriority) {
let thread = ThreadBuilder::default()
.policy(RealTime(Fifo))
.priority(priority)
.spawn(move |_| {
let mut local_executor = executor::LocalPool::new();
let spawner = local_executor.spawner();
spawner.spawn_local(future).unwrap();
local_executor.run();
[…]
thread_priority::unix::set_thread_priority_and_policy(
thread_priority::thread_native_id(),
ThreadPriority::try_from(MAIN_PRIORITY)?,
RealTime(Fifo),
[…]
let mut node = r2r::Node::create(ctx, "example", "")?;
[…]
let future = subs.for_each(move |msg: Msg| async move {
// process msg
spawn_in_thread(future, ThreadPriority::try_from(CALLBACK_PRIORITY)?);
[…]
node.spin_once(SPIN_TIMEOUT);
@@ 4 Experimental evaluation
To evaluate the proposed application structure and compare it with other alternatives, we designed a set synthetic benchmarks. These are described in Section˜4.1. Evaluation on a more complex autonomous driving application follows in Section˜4.2.
@@ 4.1 Synthetic benchmarks
To evaluate the proposed application structure and compare response times obtained with different setups, we develop a benchmark application composed of five topics (see Figure˜10). Messages are published periodically to the topics by a publisher node running in a dedicated process. The second node subscribes to the topics and executes callbacks with a specific fixed execution time. The publication periods and callback execution times are given in Table˜2 and lead to CPU utilization of 90%.
@@ 4.1.3 Implementation variants of the subscriber node
R2R-based subscriber node with futures local executor running interleaved with the spin_once function in the same thread, as in Figure˜3.
@@
The variant proposed in Section˜3, that is R2R-subscriber node with each callback running in a separate thread with real-time priority set according to the rate-monotonic (RM) priority assignment (the smaller period, the higher priority) in a futures local executor. The spin_once function is called in a loop in the main thread with the highest priority.
@@ 4.1.4 Synthetic benchmark results
To investigate the effect of RMW/DDS thread priority, we compare the end-to-end latencies of our proposed futures-rt with and without giving the RMW/DDS threads the highest priority. The results in Figure˜12 show that without setting the priority of RMW/DDS threads to be higher than the priorities of the callback threads, the callbacks fail to complete before the deadlines given by the publication periods.
@@ 4.2 Complex autonomous driving application
Our ALKS implementation can drive a real Porsche Cayenne car by connecting to its FlexRay buses and communicating with the onboard control units. However, in this paper, we evaluate the version running against the simulated vehicle in the CARLA simulator [7]. The reason is that this version contains three ROS nodes written using R2R, whereas the real car version has only two R2R-based nodes. The high-level architecture of the application is depicted in Figure˜13.
@@
The overview of threads in the nodes and their SCHED_FIFO priorities is given in Table˜3. Only the ALKS node runs callbacks in more than one thread.
@@ 4.2.1 Latency evaluation
To show that R2R is suitable for implementing practical real-time applications, we run the above-described ALKS application and trace it with LTTng. As opposed to the synthetic benchmark, here we did not restrict the threads to run on a single CPU core. From the traces, we obtain end-to-end latencies of the odometry chain (marked with red arrows in Figure˜13). This chain goes through all three R2R-based nodes in the application. Specifically, it starts at the reception of the IMU message from CARLA in the FlexRay adapter. In the associated callback, we retrieve other information from CARLA using its API and publish the obtained data to the corresponding ROS topic. The ALKS node then receives (among others) the odometry message, converts it from relative to absolute values, and stores it for later use. It also publishes the converted absolute values for use in the FlexRay visualizer. Flexray visualizer just stores the received odometry values in memory for use in other callbacks.
Figure˜14 shows histograms obtained from a trace recorded during approximately 3 minutes of execution of the ALKS application. We removed few initial samples from the trace to filter out outliers from the warm-up phase.
The histogram of measured odometry chain end-to-end latencies is shown in Figure˜14(a). In about 99% of cases, the latency is below 0.8 ms. The same figure also shows execution time (duration) of the first callback in the chain (IMU). As this callback communicates with the CARLA simulator over the network, its duration and jitter is dominated by this communication latency and this propagates further down the chain. Figure˜14(b) shows execution time histograms of other callbacks in the chain. In the ALKS node, the majority of callback executions exhibits execution time jitter of about 0.2 ms, which is caused by waiting for mutexes protecting data shared with the timer callback running in another thread. A few outliers between 0.3 and 2 ms can be caused by the fact that the experiment was executed on a machine connected to the network and running other development tools, which might cause some interference. The jitter of the callback in the FlexRay visualizer node is very low, as the node is single-threaded and does not need to use mutexes in the callbacks.
@@
Figure˜14(d) shows the execution time histogram of the timer callback in the ALKS node. The left peak represents the cases where the MPC planner was not invoked, whereas the peak around 8 ms includes the cases with planner invocations. Note that the end-to-end latency of the odometry chain is influenced by the timer callback only thourgh short critical sections for accessing shared data. Long execution time of the MPC planner has no effect on the end-to-end latency of the odometry chain. This demonstrates that R2R with the futures runtime and the structure proposed in this paper is capable of achieving deterministic end-to-end latencies of the chain of callbacks executing in different nodes.